### v2

@@ 3.2.6. Interpretability and Causal Parameters
The treatment effect parameter $\delta$ quantifies the average causal impact of interventions on target outcomes across regions. A significant negative $\delta$ (e.g., $\delta<-0.1$) indicates that policy interventions effectively reduce case counts, while $\delta\approx 0$ suggests limited intervention effectiveness. The covariate effect vector $\boldsymbol{\gamma}$ from Eq. (5) captures the marginal contribution of each epidemiological factor. For instance, $\gamma_{1}>0$ confirms that higher reproduction numbers drive increased transmission, while $\gamma_{2}<0$ validates that stricter policies reduce case growth rates. The magnitude of $|\gamma_{3}|$ reflects the relative importance of vaccination coverage in epidemic control. The spatial spillover parameter $\rho$ measures inter-regional dependency strength. Values of $\rho>0.3$ indicate strong positive spatial correlation, suggesting that neighboring regions’ epidemic patterns significantly influence local outcomes through cross-border mobility or policy coordination effects.
@@ 4. Experimental Evaluation
To assess the performance of the proposed STOAT framework on spatial-temporal causal time series (STC-TS) data, we conduct experiments using a comprehensive COVID-19 dataset. We compare STOAT against state-of-the-art probabilistic time-series forecasting methods, focusing on its ability to model inter-regional spatial dependencies and causal effects.
@@ 4.1.1. COVID-19 Dataset with Spatial and Causal Factors
  • Causal Factors: The causal factors include ICU admissions, vaccination booster rates, Stringency Index (Hale et al., 2021), and daily reproduction number, representing virus mutation. These factors constitute the covariate matrix $\mathbf{C}$. Omicron proportion data is sourced from the GISAID initiative [^1] Available at covariants.org., vaccination data from Our World in Data [^2] Available at github.com/owid/covid-19-data., Stringency Index from the Oxford Covid-19 Government Response Tracker, and reproduction rates from a public repository (Abadie, 2005) [^3] Available at github.com/crondonm/TrackingR..
@@ 4.1.2. Baseline Models
We compare STOAT with leading probabilistic time-series forecasting models implemented in the GluonTS toolkit (Alexandrov et al., 2020): DeepAR, Deep State Space Model (DSSM), DeepFactor (DF), MTSNet (Yang and Cao, 2023), and DeepVAR. Non-probabilistic methods (e.g., ARIMA, LSTM, etc.) are excluded, as they do not produce probabilistic forecasts suitable for comparison. As STOAT focuses on probabilistic forecasting to evaluate distribution accuracy, we do not compare with point-based spatio-temporal forecasting models.
  • DeepAR (Salinas et al., 2020): Utilizes a two-layer RNN to extract temporal features, mapping hidden states to distribution parameters for probabilistic predictions.
  • DSSM (Rangapuram et al., 2018): Combines a linear state-space model with RNNs, incorporating Gaussian noise for tractable likelihood-based training.
  • DF (Wang et al., 2019): Employs RNNs for global time-series representations, using Kalman filtering to generate Gaussian-based probabilistic forecasts.
  • MTSNet (Yang and Cao, 2023): Leverages global learning framework with RNNs to model multiple time series, outputting parameters for probabilistic distributions.
  • DeepVAR (Salinas et al., 2019): Extends vector autoregression with deep learning, using RNNs to model multivariate dependencies with Copula function and produce probabilistic forecasts.
@@
Time Period Metrics DeepVAR DeepAR DSSM DeepFactor MTSNet STOAT-StudentT STOAT-Gaussian STOAT-Laplace 5 Days Quantile Loss [0.1] 0.1894 0.1723 0.2015 0.1976 0.1845 0.1543 0.1421 0.1478 Quantile Loss [0.5] 0.3742 0.3421 0.3927 0.3814 0.3598 0.3476 0.3378 0.3312 Quantile Loss [0.9] 0.2653 0.2512 0.2789 0.2478 0.2634 0.2356 0.2412 0.2389 Coverage [0.1] 0.2000 0.3000 0.4000 0.1000 0.2000 0.1000 0.1000 0.1000 Coverage [0.5] 0.4000 0.6000 0.3000 0.7000 0.5000 0.4000 0.5000 0.5000 Coverage [0.9] 0.8000 0.9000 0.6000 0.8000 0.7000 0.9000 0.8000 0.9000 CRPS 0.3298 0.2954 0.3512 0.3176 0.3289 0.2897 0.2745 0.2798 Energy Score 1.2789 1.1432 1.4123 1.3214 1.2987 1.1012 1.0789 1.0876 7 Days Quantile Loss [0.1] 0.2187 0.2056 0.2312 0.2234 0.2112 0.1898 0.1789 0.1845 Quantile Loss [0.5] 0.5245 0.4876 0.5467 0.5321 0.4987 0.4723 0.4612 0.4534 Quantile Loss [0.9] 0.2812 0.2678 0.2934 0.2712 0.2545 0.2612 0.2589 0.2567 Coverage [0.1] 0.3000 0.2000 0.4000 0.1000 0.2000 0.1000 0.1000 0.1000 Coverage [0.5] 0.4000 0.6000 0.3000 0.7000 0.5000 0.4000 0.5000 0.5000 Coverage [0.9] 0.7000 0.9000 0.6000 0.8000 0.8000 0.9000 0.8000 0.9000 CRPS 0.3623 0.3478 0.3845 0.3712 0.3543 0.3176 0.3056 0.3123 Energy Score 1.4678 1.3765 1.5345 1.4567 1.4321 1.2234 1.1876 1.1987 10 Days Quantile Loss [0.1] 0.2987 0.2765 0.3123 0.2876 0.2934 0.2567 0.2456 0.2512 Quantile Loss [0.5] 0.7234 0.6987 0.7456 0.7321 0.7123 0.6654 0.6456 0.6234 Quantile Loss [0.9] 0.3789 0.3567 0.3923 0.3678 0.3545 0.3321 0.3289 0.3212 Coverage [0.1] 0.3000 0.2000 0.4000 0.0000 0.2000 0.1000 0.1000 0.1000 Coverage [0.5] 0.3000 0.7000 0.2000 0.6000 0.5000 0.4000 0.4000 0.4000 Coverage [0.9] 0.6000 0.7000 0.5000 0.7000 0.8000 0.9000 0.8000 0.9000 CRPS 0.4876 0.4654 0.5123 0.4987 0.4765 0.4321 0.4123 0.4234 Energy Score 1.8765 1.7898 1.9456 1.8765 1.8321 1.5432 1.4987 1.5123
Table 1. Comparison of probabilistic forecasting performance across models for 5-day, 7-day, and 10-day horizons on COVID-19 case data, averaged over six countries. Metrics include Quantile Loss at 0.1, 0.5, and 0.9 quantiles, Coverage at corresponding quantiles, Continuous Ranked Probability Score (CRPS), and Energy Score. The proposed STOAT variants (Student’s-$t$, Gaussian, Laplace) are compared against baseline models (DeepVAR, DeepAR, DSSM, DeepFactor, MTSNet). Bold values indicate the best performance for each metric and time period.
@@ 4.2. Comparison Results
Figure 3. Visualization of 10-day probabilistic forecasts for COVID-19 case data during peak periods in Canada (left) and Italy (right) using the STOAT model. The yellow, green, and red lines represent the predictive distributions based on Student’s-$t$, Gaussian, and Laplace distributions, respectively.
@@
4.3. Ablation Study
Figure 6. Absolute performance of the STOAT model and its ablation variants across three distribution choices (Student’s-$t$, Gaussian, Laplace) and three forecast horizons (5, 7, 10 days). For each distribution, the full model (“Full”) is compared with its “NoSpatial” and “NoFactors” ablation variants, reporting both CRPS (bars, left axis) and Energy Score (lines, right axis). The results show that removing either spatial causal adjustment or causal factor integration generally leads to degraded predictive performance, confirming the contribution of both components to overall model accuracy.
To evaluate the contributions of spatial causal inference and causal factor integration within STOAT, we perform ablation experiments on COVID-19 case data, averaged over peak periods in six countries. Across three probabilistic output distributions (Laplace, Gaussian, Student’s-$t$ distribution), we compare the full model against ablated variants that remove either the spatial relation matrix or the set of causal covariates.
As shown in Table 2, the complete STOAT model consistently achieves the best overall performance in terms of CRPS and Energy Score across all forecast horizons. For instance, STOAT-Gaussian achieves a CRPS of 0.2745 and an Energy Score of 1.0789 on the 5-day horizon, outperforming both -NoSpatial and -NoFactors variants. However, we observe that in some cases, ablated models yield lower errors on individual metrics. For example, STOAT-StudentT-NoSpatial attains a lower Quantile Loss [0.1] (0.1512) than the full Student’s-$t$ variant (0.1543), and STOAT-Gaussian-NoFactors achieves a lower Quantile Loss [0.1] (0.1400) compared to the full Gaussian model (0.1421) on the same horizon. This phenomenon is most pronounced on short-term forecasts, where reduced model complexity or incidental alignment with test-set fluctuations can occasionally result in marginally better pointwise quantile performance.
Nevertheless, as the prediction horizon increases, the benefits of joint spatial and causal modeling become more apparent. For example, on the 10-day forecast, the full STOAT-Gaussian model yields a CRPS of 0.4123, compared to 0.4204 and 0.4229 for the -NoSpatial and -NoFactors ablations, respectively, with a similar trend for the Laplace and Student’s-$t$ settings. These results highlight that while ablated models may occasionally surpass the full model on isolated quantile losses, the overall probabilistic calibration and sharpness—as captured by proper scoring rules—remain superior for the complete spatial-causal STOAT. This underscores the importance of incorporating both spatial dependencies and causally relevant covariates for robust spatial-temporal probabilistic forecasting.
@@ 4.4. Conclusion
5. Acknowledgments
This research is supported by the Australian Commonwealth Scientific and Industrial Research Organisation (CSIRO) and the the United States National Science Foundation under Grants No. 2302968, No. 2302969, and No. 2302970, titled ”Collaborative Research: NSF-CSIRO: HCC: Small: Understanding Bias in AI Models for the Prediction of Infectious Disease Spread”.
@@ A.2. Metrics
  • Continuous Ranked Probability Score (CRPS) (Gneiting et al., 2005): Quantifies the discrepancy between predicted and empirical cumulative distribution functions:
@@
    $Q_{i,t}^{(\tau)}$ =$=\begin{cases}(1-\tau)|q_{i,t}^{(\tau)}-x_{i,t}|&\text{if }q_{i,t}^{(\tau)}>x_{i,t},\\ \tau|q_{i,t}^{(\tau)}-x_{i,t}|&\text{otherwise},\end{cases}$
@@
  • Energy Score (Jordan et al., 2019): Evaluates multivariate probabilistic forecast accuracy, sensitive to inter-variable correlations:

### v3

@@ 3.2.6. Interpretability and Causal Parameters
The treatment effect parameter $\delta$ quantifies the average causal impact of interventions on target outcomes across regions. A significant negative $\delta$ (e.g., $\delta<-0.1$) indicates that policy interventions effectively reduce case counts, while $\delta\approx 0$ suggests limited intervention effectiveness.
The covariate effect vector $\boldsymbol{\gamma}$ from Eq. (5) captures the marginal contribution of each epidemiological factor. For instance, $\gamma_{1}>0$ confirms that higher reproduction numbers drive increased transmission, while $\gamma_{2}<0$ validates that stricter policies reduce case growth rates. The magnitude of $|\gamma_{3}|$ reflects the relative importance of vaccination coverage in epidemic control.
The spatial spillover parameter $\rho$ measures inter-regional dependency strength. Values of $\rho>0.3$ indicate strong positive spatial correlation, suggesting that neighboring regions’ epidemic patterns significantly influence local outcomes through cross-border mobility or policy coordination effects.
@@ 4. Experimental Evaluation
To assess the performance of the proposed STOAT framework on spatial-temporal causal multivariate time series (STC-MTS) data, we conduct experiments using a comprehensive COVID-19 dataset. We compare STOAT against state-of-the-art probabilistic time-series forecasting methods, focusing on its ability to model inter-regional spatial dependencies and causal effects.
@@ 4.1.1. COVID-19 Dataset with Spatial and Causal Factors
  • Causal Factors: The causal factors include ICU admissions, vaccination booster rates, Stringency Index (hale2021global), and daily reproduction number, representing virus mutation. These factors constitute the covariate matrix $\mathbf{C}$. Omicron proportion data is sourced from the GISAID initiative [^1] Available at covariants.org., vaccination data from Our World in Data [^2] Available at github.com/owid/covid-19-data., Stringency Index from the Oxford Covid-19 Government Response Tracker, and reproduction rates from a public repository (abadie2005semiparametric) [^3] Available at github.com/crondonm/TrackingR..
@@ 4.1.2. Baseline Models
We compare STOAT with leading probabilistic time-series forecasting models implemented in the GluonTS toolkit (alexandrov2020gluonts): DeepAR, Deep State Space Model (DSSM), DeepFactor (DF), MTSNet (yang2023mtsnet), and DeepVAR. Non-probabilistic methods (e.g., ARIMA, LSTM, etc.) are excluded, as they do not produce probabilistic forecasts suitable for comparison. As STOAT focuses on probabilistic forecasting to evaluate distribution accuracy, we do not compare with point-based spatio-temporal forecasting models.
  • DeepAR (salinas2020deepar): Utilizes a two-layer RNN to extract temporal features, mapping hidden states to distribution parameters for probabilistic predictions.
  • DSSM (rangapuram2018deep): Combines a linear state-space model with RNNs, incorporating Gaussian noise for tractable likelihood-based training.
  • DF (wang2019deep): Employs RNNs for global time-series representations, using Kalman filtering to generate Gaussian-based probabilistic forecasts.
  • MTSNet (yang2023mtsnet): Leverages global learning framework with RNNs to model multiple time series, outputting parameters for probabilistic distributions.
  • DeepVAR (salinas2019high): Extends vector autoregression with deep learning, using RNNs to model multivariate dependencies with Copula function and produce probabilistic forecasts.
@@
Time Period Metrics DeepVAR DeepAR DSSM DeepFactor MTSNet STOAT-Student-T STOAT-Gaussian STOAT-Laplace 5 Days Quantile Loss [0.1] 0.1894 0.1723 0.2015 0.1976 0.1845 0.1543 0.1421 0.1478 Quantile Loss [0.5] 0.3742 0.3421 0.3927 0.3814 0.3598 0.3476 0.3378 0.3312 Quantile Loss [0.9] 0.2653 0.2512 0.2789 0.2478 0.2634 0.2356 0.2412 0.2389 Coverage [0.1] 0.2000 0.3000 0.4000 0.1000 0.2000 0.1000 0.1000 0.1000 Coverage [0.5] 0.4000 0.6000 0.3000 0.7000 0.5000 0.4000 0.5000 0.5000 Coverage [0.9] 0.8000 0.9000 0.6000 0.8000 0.7000 0.9000 0.8000 0.9000 CRPS 0.3298 0.2954 0.3512 0.3176 0.3289 0.2897 0.2745 0.2798 Energy Score 1.2789 1.1432 1.4123 1.3214 1.2987 1.1012 1.0789 1.0876 7 Days Quantile Loss [0.1] 0.2187 0.2056 0.2312 0.2234 0.2112 0.1898 0.1789 0.1845 Quantile Loss [0.5] 0.5245 0.4876 0.5467 0.5321 0.4987 0.4723 0.4612 0.4534 Quantile Loss [0.9] 0.2812 0.2678 0.2934 0.2712 0.2545 0.2612 0.2589 0.2567 Coverage [0.1] 0.3000 0.2000 0.4000 0.1000 0.2000 0.1000 0.1000 0.1000 Coverage [0.5] 0.4000 0.6000 0.3000 0.7000 0.5000 0.4000 0.5000 0.5000 Coverage [0.9] 0.7000 0.9000 0.6000 0.8000 0.8000 0.9000 0.8000 0.9000 CRPS 0.3623 0.3478 0.3845 0.3712 0.3543 0.3176 0.3056 0.3123 Energy Score 1.4678 1.3765 1.5345 1.4567 1.4321 1.2234 1.1876 1.1987 10 Days Quantile Loss [0.1] 0.2987 0.2765 0.3123 0.2876 0.2934 0.2567 0.2456 0.2512 Quantile Loss [0.5] 0.7234 0.6987 0.7456 0.7321 0.7123 0.6654 0.6456 0.6234 Quantile Loss [0.9] 0.3789 0.3567 0.3923 0.3678 0.3545 0.3321 0.3289 0.3212 Coverage [0.1] 0.3000 0.2000 0.4000 0.0000 0.2000 0.1000 0.1000 0.1000 Coverage [0.5] 0.3000 0.7000 0.2000 0.6000 0.5000 0.4000 0.4000 0.4000 Coverage [0.9] 0.6000 0.7000 0.5000 0.7000 0.8000 0.9000 0.8000 0.9000 CRPS 0.4876 0.4654 0.5123 0.4987 0.4765 0.4321 0.4123 0.4234 Energy Score 1.8765 1.7898 1.9456 1.8765 1.8321 1.5432 1.4987 1.5123
Table 1. Comparison of probabilistic forecasting performance across models for 5-day, 7-day, and 10-day horizons on COVID-19 case data, averaged over six countries. Metrics include Quantile Loss at 0.1, 0.5, and 0.9 quantiles, Coverage at corresponding quantiles, Continuous Ranked Probability Score (CRPS), and Energy Score. The proposed STOAT variants (Student-$t$, Gaussian, Laplace) are compared against baseline models (DeepVAR, DeepAR, DSSM, DeepFactor, MTSNet). Bold values indicate the best performance for each metric and time period.
@@ 4.2. Comparison Results
Figure 3. Visualization of 10-day probabilistic forecasts for COVID-19 case data during peak periods in Canada (left) and Italy (right) using the STOAT model. The yellow, green, and red lines represent the predictive distributions based on Student-t, Gaussian, and Laplace distributions, respectively.
@@
4.3. Ablation Study
To evaluate the contributions of spatial causal inference and causal factor integration within STOAT, we perform ablation experiments on COVID-19 case data, averaged over peak periods in six countries. Across three probabilistic output distributions (Laplace, Gaussian, Student-$t$), we compare the full model against ablated variants that remove either the spatial relation matrix or the set of causal covariates.
As shown in Table 2, the complete STOAT model consistently achieves the best overall performance in terms of CRPS and Energy Score across all forecast horizons. For instance, STOAT-Gaussian achieves a CRPS of 0.2745 and an Energy Score of 1.0789 on the 5-day horizon, outperforming both -NoSpatial and -NoFactors variants. However, we observe that in some cases, ablated models yield lower errors on individual metrics. For example, STOAT-StudentT-NoSpatial attains a lower Quantile Loss [0.1] (0.1512) than the full Student-$t$ variant (0.1543), and STOAT-Gaussian-NoFactors achieves a lower Quantile Loss [0.1] (0.1400) compared to the full Gaussian model (0.1421) on the same horizon. This phenomenon is most pronounced on short-term forecasts, where reduced model complexity or incidental alignment with test-set fluctuations can occasionally result in marginally better pointwise quantile performance.
Nevertheless, as the prediction horizon increases, the benefits of joint spatial and causal modeling become more apparent. For example, on the 10-day forecast, the full STOAT-Gaussian model yields a CRPS of 0.4123, compared to 0.4204 and 0.4229 for the -NoSpatial and -NoFactors ablations, respectively, with a similar trend for the Laplace and Student-$t$ settings. These results highlight that while ablated models may occasionally surpass the full model on isolated quantile losses, the overall probabilistic calibration and sharpness—as captured by proper scoring rules—remain superior for the complete spatial-causal STOAT. This underscores the importance of incorporating both spatial dependencies and causally relevant covariates for robust spatial-temporal probabilistic forecasting.
@@ 4.3. Ablation Study
Figure 6. Absolute performance of the STOAT model and its ablation variants across three distribution choices (Student-T, Gaussian, Laplace) and three forecast horizons (5, 7, 10 days). For each distribution, the full model (“Full”) is compared with its “NoSpatial” and “NoFactors” ablation variants, reporting both CRPS (bars, left axis) and Energy Score (lines, right axis). The results show that removing either spatial causal adjustment or causal factor integration generally leads to degraded predictive performance, confirming the contribution of both components to overall model accuracy.
@@ A.2. Metrics
  • Continuous Ranked Probability Score (CRPS) (gneiting2005calibrated): Quantifies the discrepancy between predicted and empirical cumulative distribution functions:
@@
[…]
@@
  • Energy Score (jordan2019evaluating): Evaluates multivariate probabilistic forecast accuracy, sensitive to inter-variable correlations: